OpenStudy (anonymous):

he temperature was 65 degrees at daybreak. Then it dropped 2 degrees per hour until dusk. If x is the number of hours since daybreak, this equation represents the temperature during the day: y = -2x + 65 OR y = 65 - 2x. Use the equation to predict the temperature 6 hours later.

OpenStudy (anonymous):

x is the number of hours after daybreak y is his temperature after x hours so , y = 65 - 2(6) since x = 6 hrs = 53

OpenStudy (anonymous):

Thank you. Can you help me on this one... What is the slope in the equation: y = 6x - 300? A. 300 B. 6 C.-6 D. -300

OpenStudy (anonymous):

equation of a straight line: y = mx + c m is the gradient so whatever is paired with x is the gradient hence the answer is 6 c is the y-intercept so the graph cuts the vertical axis at y = -300

OpenStudy (anonymous):

Okay, thank you.

OpenStudy (anonymous):

You drive a car that holds 15 gallons of gasoline. Your car uses 0.03 gal per mi. Which equation best represents this information? A. y = 0.03x + 15 B. y = 0.03x - 15 C. y = -0.03x + 15 D. y = -0.03x - 15

OpenStudy (anonymous):

C or D

OpenStudy (anonymous):

Right?

OpenStudy (anonymous):

Adam goes to a skating rink on Tuesdays for $8.50. While there, he pays $1.50 for each item he buys at the concession stand. Which equation best represents this information?